OpenStudy (anonymous):

the no. of triplets of a,b,c for which the system of equations ax+by=2a-b & (c+1)x+cy=10-a+3b has infinetely many solutions and , x=1 and y=3 is one of the solutions is, (A) exactly one (B)exactly two (C)exactly three (D)infinitly many. pLz explain.

OpenStudy (anonymous):

can u explain it

OpenStudy (anonymous):

u can't xplain it

OpenStudy (anonymous):

i think ur typing speed is very slow so u hav 2 work on that.

OpenStudy (anonymous):

So we know that \[a+3b=2-3\implies a=-3b-1\]and\[4c=10+6b \implies c=\frac{5+3b}{2}\]Seems if we pick a value for b, we can get values for a and c that make it work for point (1,3). This indicates infinitely many triples (a,b,c) that can solve for that point. If the system is dependent, \[ka=c+1, kb=c,~and~k(2a-b)=10-3a+b\]

OpenStudy (anonymous):

After a little substitution,\[k(-3b-1)=\frac{3b+7}{2},~kb=\frac{5+3b}{2},~k(-7b-2)=10+10b+3\]which means we need to solve\[k=\frac{3b+7}{2(-3b-1)}=\frac{5+3b}{2b}=\frac{13+10b}{-7b-2}\]

OpenStudy (anonymous):

After a little more work,\[6b^2+14b=-18b^2-36b-10~and~20b^2+26b=-21b^2-41b-10\]simplify a little to get\[17b^2-17b-20=0\]This has two roots, so we expect two solutions. This is sort of circuitous, and there was enough algebra involved that I can in no way guarantee my calculations, but these posts should give you enough structure to investigate the problem.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

After looking a little more, I think the last polynomial is\[17b^2-17b=0\]which would indicate a single solution (we've had b in the denominator, so it can't be zero). Like I said before, there were quite a few tedious calculations, and I'm not noted for my flawless arithmetic.